# Reinforcement Learning in Vision: A Survey

### v1

@@ 5.3 RL for Thinking with Vision
Recent works for visual planning, such as Chain-of-Focus (Zhang et al., 2025f) and Openthinkimg (Su et al., 2025b) all treat the picture as an external workspace: the agent may crop, sketch, highlight or insert visual tokens before emitting the next language token. While early prototypes rely on supervised heuristics for these spatial actions, moving to reinforcement learning exposes four open problems. (i) Action-space design. Cropping or doodling is naturally continuous ($x,y,w,h,\dots$) yet RL libraries and GPU memories favor small discrete sets. Hybrid schemes that learn a differentiable proposal policy and then refine coordinates via policy gradient fine-tuning, as hinted by BRPO (Chu et al., 2025) and VRAG-RL (Wang et al., 2025e), remain largely unexplored. (ii) Credit assignment. Most benchmarks only reward the final task success (e.g., answer correctness in VILASR (Wu et al., 2025a)); the whole visual chain-of-thought shares a single sparse scalar. Future work should mine step-wise proxy rewards, e.g., CLIP similarity increase after a crop, or entropy drop in a learned belief state—to enable bootstrapped or hierarchical RL. (iii) Data efficiency. Sketching or inserting patches triggers extra forward passes through the vision encoder, making naive on-policy RL prohibitively expensive. Relabeling (DeepEyes (Zheng et al., 2025)) and model-based imagination (Pixel Reasoner (Su et al., 2025a)) point to sample-efficient alternatives, but principled replay and uncertainty-aware planners for visual actions are still missing. Therefore, future directions include: learning structured visual skills (crop, zoom, draw) via skill-prior RL; devising cross-modal reward shaping that scores each edit by how much it simplifies the remaining reasoning; and curating benchmarks whose metrics expose not just final accuracy.
@@ 5.4 Reward Model Design for Visual Generation
A central obstacle for reinforcement-learning–based visual generation is the lack of a scalable and faithful reward function. Widely used handcrafted metrics such as FID (Heusel et al., 2017) offer a convenient numerical signal, yet correlate only weakly with human judgments of aesthetics, semantic fidelity, or temporal coherence, especially when the task extends beyond single-frame images. Recent learned critics, such as ImageReward (Xu et al., 2023) and HPS (Wu et al., 2023b) for images, and VideoReward (Liu et al., 2025a) for videos to bridge this gap by training on pairwise human-preference data, but each model targets a narrow modality and captures only a slice of perceptual quality (e.g., prompt alignment or visual appeal). As a result, policies optimized with PPO or GRPO often exploit loopholes in a single scalar signal, producing high-contrast artifacts, repetitive textures, or physically implausible motion that “game” the critic without improving real user satisfaction. The challenge, therefore, is to design reward models that (i) integrate complementary low-level signals (consistency, physics, geometry) with high-level human preferences, (ii) generalize across images, video and 3-D scenes, and (iii) remain robust against reward hacking while being cheap enough to update continually as user tastes evolve.

### v2

@@ 5.3 RL for Thinking with Vision
Recent works for visual planning, such as Chain-of-Focus (Zhang et al., 2025g) and Openthinkimg (Su et al., 2025b) all treat the picture as an external workspace: the agent may crop, sketch, highlight or insert visual tokens before emitting the next language token. While early prototypes rely on supervised heuristics for these spatial actions, moving to reinforcement learning exposes four open problems. (i) Action-space design. Cropping or doodling is naturally continuous ($x,y,w,h,\dots$) yet RL libraries and GPU memories favor small discrete sets. Hybrid schemes that learn a differentiable proposal policy and then refine coordinates via policy gradient fine-tuning, as hinted by BRPO (Chu et al., 2025) and VRAG-RL (Wang et al., 2025f), remain largely unexplored. (ii) Credit assignment. Most benchmarks only reward the final task success (e.g., answer correctness in VILASR (Wu et al., 2025a)); the whole visual chain-of-thought shares a single sparse scalar. Future work should mine step-wise proxy rewards, e.g., CLIP similarity increase after a crop, or entropy drop in a learned belief state—to enable bootstrapped or hierarchical RL. (iii) Data efficiency. Sketching or inserting patches triggers extra forward passes through the vision encoder, making naive on-policy RL prohibitively expensive. Relabeling (DeepEyes (Zheng et al., 2025)) and model-based imagination (Pixel Reasoner (Su et al., 2025a)) point to sample-efficient alternatives, but principled replay and uncertainty-aware planners for visual actions are still missing. Therefore, future directions include: learning structured visual skills (crop, zoom, draw) via skill-prior RL; devising cross-modal reward shaping that scores each edit by how much it simplifies the remaining reasoning; and curating benchmarks whose metrics expose not just final accuracy.
@@ 5.4 Reward Model Design for Visual Generation
A central obstacle for reinforcement-learning–based visual generation is the lack of a scalable and faithful reward function. Widely used handcrafted metrics such as FID (Heusel et al., 2017) offer a convenient numerical signal, yet correlate only weakly with human judgments of aesthetics, semantic fidelity, or temporal coherence, especially when the task extends beyond single-frame images. Recent learned critics, such as ImageReward (Xu et al., 2023) and HPS (Wu et al., 2023b) for images, and VideoReward (Liu et al., 2025b) for videos to bridge this gap by training on pairwise human-preference data, but each model targets a narrow modality and captures only a slice of perceptual quality (e.g., prompt alignment or visual appeal). As a result, policies optimized with PPO or GRPO often exploit loopholes in a single scalar signal, producing high-contrast artifacts, repetitive textures, or physically implausible motion that “game” the critic without improving real user satisfaction. The challenge, therefore, is to design reward models that (i) integrate complementary low-level signals (consistency, physics, geometry) with high-level human preferences, (ii) generalize across images, video and 3-D scenes, and (iii) remain robust against reward hacking while being cheap enough to update continually as user tastes evolve.